OpenStudy (anonymous):

Please Help Me. There will be a medal for those who will answer this question. Describe the ways of recognizing a quadratic function. In: a. Graph b. Equation c. Table Values Good Luck. And Thanks.

OpenStudy (anonymous):

Its a table value

OpenStudy (welshfella):

The graph will be a parabola looks like

OpenStudy (welshfella):

- which might also open downwards or sideways The curve is symmetrical about a line passing through its vertex

OpenStudy (welshfella):

b the degree of a quadratic equation will be 2. Meaning in will contain a squared variable as its highest exponent. eg x^2 = 0 , x^2 - x - 10 = 0 3y^2 + 4y - 7 = 0

OpenStudy (welshfella):

c Table the values of f(x) will increase/decrease by an increasing amount and the increases will be in arithmetic progression. (I'm not sure if this is always the case)

OpenStudy (welshfella):

an example would be f(x) = x^2 + 2 x f(x) -- --- 1 3 2 6 3 11 4 18 as you can see the differences in value of f(x) are 3 , 5 , 7 which is an AS.